A plane flies 720 mi against a steady 30-mi/h headwind and then returns to the same point with the wind. If the entire trip takes 10h, what is the plane's speed still in the air?
Answer by checkley71(8403)

720/(X+30)+720/(X-30)=10 OR [720(X-30)+720(X+30)]/(X~2-900)=10 OR
(720X-21600+720X-21600)/(X~2-900)=10 OR 10(X~2-900)=1440X OR X~2-144X-900=0 OR
(X-150)(X+6)=0 OR X=150 & X=-6 THEREFORE THE SPEED OF THE PLANE IS 150MPH